OpenStudy (anonymous):

If a chord of circle with radius 5 cm subtends a right angle at the centre , then length of chord is

OpenStudy (anonymous):

Draw the Pic First :

OpenStudy (anonymous):



OpenStudy (anonymous):

Use pythagorus Theorem....

OpenStudy (anonymous):

\[Chord = \sqrt{5^2+5^2}=\sqrt{50}=5\sqrt{2}\]

OpenStudy (anonymous):

if it subtend a right a right angle it subtends 90 degrees therefore lenght of chord=pi rx/180 =pi(5)90/180 = 5 (pi)/2=7.854 cm.....ans